# Nonlinear Multi-objective Reinforcement Learning with Provable Guarantees

### v2

@@ 5 The RA-Value Iteration Algorithm
Theorem 12 (Optimality Guarantee of RA-Value Iteration).
For a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $|\mathcal{S}|$, and $|\mathcal{A}|$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$, and exponential in the number of objectives $d$, and $A$ will output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
Proof.
Note that in particular, Lemma 11 implies that $V(s,\mathbf{0},T)\geq V^{*}(s,\mathbf{0},T)-T\epsilon$. This implies that for a finite time horizon $T$ and a desired approximation error $\epsilon$, it suffices to choose $\alpha_{T\epsilon}=\delta_{T\epsilon}/d$ where $\delta_{T\epsilon}$ comes from the smoothness of the welfare function. One then uses RA-Value Iteration. ∎
6 Removing Knowledge of the Model
We have shown that RA-value Iteration can find an approximately optimal policy in polynomial time (for a constant number of objectives) given access to a model of the environment. Next, we describe how to learn the model by extending the classical E3 algorithm (Kearns and Singh 2002) to the multi-objective case. We call the resulting algorithm Reward-Aware Explicit Explore or Exploit (RA-E3).
A learning algorithm cannot recover from low return during the early part of learning due to the exponentially decaying effects of the discounting factor. Therefore, our goal for a learning algorithm is to efficiently compute a policy while exploring so that, starting from the current state with 0 initial accumulated reward and time steps remaining reset to $T$, that policy has expected welfare within $\epsilon$ of the best possible beginning from that state.
For ease of exposition, we will describe the algorithm under the assumption that we know the optimal return $V^{*}(s,\mathbf{0},T)$. This assumption can be further removed via an exploration-biased approach during the offline policy computation stage similar to the original E3 algorithm (Kearns and Singh 2002). We can now state the main result yielded by the RA-E3 algorithm.
Theorem 13 (RA-E3).
Let $V^{*}(s,0,T)$ denote the value function for the policy with the optimal expected welfare in the MOMDP $M$ starting at state $s$, with $\mathbf{0}\in{\mathbb{R}}^{d}$ accumulated reward and $T$ timesteps remaining. Then for a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, and $V^{*}(s,\mathbf{0},T)$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $1/\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$ and exponential in the number of objectives $d$, and with probability at least $1-\beta$, $A$ will halt in a state $s$, and output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}_{M}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
The algorithm is mostly a straightforward extension of the E3 algorithm expanded to incorporate a multi-objective (state, action)-pair value function. Most proofs are identical apart from this or the addition of extra $|\mathcal{A}|$ runtime factors. The notable exception are the exploitation subroutines which we have introduced already. Here we sketch the argument by outlining the entire RA-E3 algorithm and emphasizing the role of our RA-Value Iteration optimization algorithm. We provide additional details comparing our analysis with that of (Kearns and Singh 2002) in the appendix.
Introducing RA-E3
@@ 6 Removing Knowledge of the Model
The extension works as follows. The algorithm starts off by doing balanced wandering (Kearns and Singh 2002): when encountering a new state, the algorithm selects a random action. However, when revisiting a previously visited state, it chooses the least attempted action from that state, resolving ties by random action selection. At each state-action pair $(s,a)$ it tries, the algorithm stores the reward $\mathbf{R}(s,a)$ received and an estimate of the transition probabilities $Pr(s^{\prime}\mid s,a)$ derived from the empirical distribution of next states reached during balanced wandering.
@@
Next, we introduce the notion of a known state (Kearns and Singh 2002), which refers to a state that the algorithm has explored to the extent that the estimated transition probabilities for any action from that state closely approximate their actual values. Denote the number of times a state needs to be visited as $m_{known}$. We will specify the value later in our runtime characterization.
@@
States are thus categorized into three groups: known states, which the algorithm has extensively visited and obtained reliable transition statistics; states that have been visited before but remain unknown due to limited trials and therefore unreliable data; and states that have not been explored at all. By the Pigeonhole Principle, accurate statistics will eventually accumulate in some states over time, leading to their becoming known. Let $S$ be the set of currently known states, the algorithm can build the current known-state MOMDP $M_{S}$ that is naturally induced on $S$ by the full MOMDP $M$ with all “unknown” states merged into a single absorbing state. Although the algorithm cannot access $M_{S}$ directly, it will have an approximation $\hat{M}_{S}$ by the definition of the known states. By the simulation lemma (Kearns and Singh 2002), $\hat{M}_{S}$ will be an accurate model in the sense that the expected $T$-step welfare of any policy in $\hat{M}_{S}$ is close to its expected $T$-step return in $M_{S}$. (Here $T$ is the horizon time.) Hence, at any timestep, $\hat{M}_{S}$ functions as an incomplete representation of $M$, specifically focusing on the part of $M$ that the algorithm possesses a strong understanding of.
@@
This is where we insert RA-Value Iteration. The algorithm performs the two off-line optimal policy computations; i) first on $\hat{M}_{S}$ using RA-Value Iteration to compute an exploitation policy that yields an approximately optimal welfare and ii) second performing traditional value iteration on $\hat{M}^{\prime}_{S}$, which has the same transition probabilities as $\hat{M}_{S}$, but different payoffs: in $\hat{M}^{\prime}_{S}$, the absorbing state (representing “unknown” states) has scalar reward $1$ and all other states have scalar reward 00. The optimal policy in $\hat{M}^{\prime}_{S}$ simply exits the known model as rapidly as possible, rewarding exploration.
@@
By the explore or exploit lemma (Kearns and Singh 2002), the algorithm is guaranteed to either output a policy with approximately optimal return in $M$, or to improve the statistics at an unknown state. Again by the Pigeonhole Principle, a new state becomes known after the latter case occurs for some finite number of times, and thus the algorithm is always making progress. In the worst case, the algorithm builds a model of the entire MOMDP $M$. Having described the elements, we now outline the entire extended algorithm, where the notations are consistent with (Kearns and Singh 2002).
RA-E3 Algorithm:
@@
  • (Discovery of New Known States) Any time a state $s$ has been visited $m_{known}$ (Kearns and Singh 2002) times during balanced wandering, it enters the known set $S$, and no longer participates in balanced wandering.
@@
    • (Attempted Exploitation) Use RA-Value Iteration algorithm to compute an $\epsilon/2$-optimal policy on $\hat{M}_{S}$. If the resulting exploitation policy $\hat{\pi}$ achieves return from $s$ in $\hat{M}_{S}$ that is at least $V^{*}(s,\mathbf{0},T)-\epsilon/2$, the algorithm halts and outputs $\hat{\pi}$.
@@
This concludes the description of the algorithm. The failure parameter $\beta$ in the statement of the main theorem can be handled similar to the original algorithm, giving a maximum number of attempted explorations of
where $G^{T}_{\max}\leq W(\mathbf{T})$ denotes the maximum possible $T$-step welfare (recall $\mathbf{R}(s,a)\in[0,1]^{d}~{}\forall s,a$).
$m_{known}=O(((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})/\epsilon)^{4}|\mathcal{A}|\log(1/\beta))$ by the same analysis as (Kearns and Singh 2002). Note that this bound differs from the original one by a factor of $|\mathcal{A}|^{5}$. This is because i) we no longer assume the number of actions be a constant as in the original assumption, which accounts for $|\mathcal{A}|$ outside the quartic, and ii) the optimal policy $\pi$ is no longer stationary with respect to the current state, which accounts for the $|\mathcal{A}|$ inside the quartic.
Thus, the total computation time is bounded by $O\left(|\mathcal{S}||\mathcal{A}|(T/\alpha_{T\epsilon/2})^{d}\right)$ (the time required for RA-Value Iteration during off-line computations with $\epsilon/2$ precision by Lemma 11) times $T$ times the maximum number of attempted explorations, giving
$O((|\mathcal{S}|^{2}|\mathcal{A}|TG^{T}_{\max}/\epsilon)(T/\alpha_{T\epsilon/2})^{d}\log(|\mathcal{S}|/\beta)m_{known}).$
This concludes the sketch of the RA-E3 theorem.
Appendix
Appendix A Example and Practical Remarks
The RA-E3 algorithm has many parameters and terms in the analysis. In this section, we provide a concrete example of a particular welfare function, the setting of relevant parameters, and a derivation of a simplified runtime analysis.
Smoothed Proportional Fairness Welfare Function.
We will use the smoothed proportional fairness objective for our example: $W_{SPF}(\mathbf{x})=\sum_{i=1}^{d}\ln(x_{i}+1)$. The sum of logarithms is known as the proportional fairness objective - note that it is simply the log-transform of the Nash Social Welfare (or geometric mean) with better numerical stability. The addition of 1 inside the log is a smoothing factor intended to stabilize behavior near 0.
Calculating $\delta_{\epsilon}$.
Uniform continuity is a weaker assumption than the more common form of $L$-Lipshitz continuity for some constant $L$. Note that for an $L$-Lipshitz function (equivalently, for a differentiable function with the magnitude of the gradient bounded everywhere on the domain by $L$), then the correct value of $\delta_{\epsilon}$ is just $\epsilon/L$, where $\epsilon$ is the desired approximation factor of the algorithm.
By taking the gradient, we can see that the $W_{SPF}(\mathbf{x})$ is Lipshitz-continuous with Lipshitz constant $d$ for nonnegative $\mathbf{x}$, where $d$ is the dimension of vector reward, so $\delta_{\epsilon}=\epsilon/d$.
Why The Discount Factor $\gamma$ Matters: Calculating The Horizon Time $T$ and maximum return $G^{T}_{max}$.
For tasks with a finite time Horizon, $T$ is a modeling choice. Otherwise, for a discounting factor $\gamma<1$, $T$ must be large enough to ensure that planning beyond the horizon time is inconsequential in terms of total discounted reward. By Lemma 8, we need
which for our running example, can be simplified to
$T\geq\left(\frac{1}{1-\gamma}\right)\log\left(\frac{d}{\epsilon(1-\gamma)}\right).$
Recall that $G^{T}_{max}$ is the maximum $T$-step welfare possible, trivially bounded by the welfare of $T$ times the identity vector (recall our rewards are normalized to at most 1 in every dimension). So in our running example,
$G^{T}_{max}\leq d\ln(1+T).$
Setting The Discretization Parameter $\alpha_{T\epsilon}$ in The RA-Value Iteration Algorithm.
Recall from Theorem 12 that we should set the discretization parameter in the RA-Value Iteration Algorithm to $\alpha_{T\epsilon}=\delta_{T\epsilon}/d$ where $\delta_{T\epsilon}$ is sufficient to drive the Lemma 11 bound to $\epsilon$, that is, it should be
$\alpha_{T\epsilon}=\frac{\delta_{\epsilon}}{T\cdot d}.$
In our running example, since $\delta_{\epsilon}=\frac{\epsilon}{d}$ so we should set
$\alpha_{T\epsilon}=\frac{\epsilon}{T\cdot d^{2}}.$
Note that concave welfare functions exhibit better smoothness by moving further away from the origin. A further optimization of the RA-Value Iteration algorithm can exploit this property by “coarsening” the grid further away from the origin, i.e. reducing grid density in regions more distant from the origin, thereby achieving equivalent approximation quality with better efficiency.
Runtime of RA-Value Iteration using $\alpha_{T\epsilon}$.
The runtime of RA-Value Iteration on our running example would be
$O\left(|\mathcal{S}||\mathcal{A}|(T/\alpha_{T\epsilon})^{d}\right)$ $=O\left(|\mathcal{S}||\mathcal{A}|\left(\frac{T^{2}\cdot d^{2}}{\epsilon}\right)^{d}\right).$
It may be helpful to note that, if one takes $|\mathcal{A}|$, the maximum number of actions available from a given state, the discount factor $\gamma$, and the dimension of reward $d$ to be small constants close to 1, then the asymptotic dependence of the runtime on the two fundamental parameters of the size of the state space $|\mathcal{S}|$ and the desired approximation $\epsilon$ is
$O\left(|\mathcal{S}|\left(\frac{1}{\epsilon}\log^{2}(1/\epsilon)\right)^{d}\right)$
for our running example.
Model Learning and and $m_{known}$.
The overall RA-E3 algorithm must learn a model during exploration, unlike the RA-Value Iteration that takes a model as input. $m_{known}$ is the number of times we must visit a state in order to “know” its statistics (especially transition probabilities) sufficiently well. In the worst case bound we give that
$m_{known}=O(((|\mathcal{S}||\mathcal{A}|TG^{T}_{\max})/\epsilon)^{4}|\mathcal{A}|\log(1/\beta))$
but this is extremely pessimistic, as it assumes that every state may transition to every other possible state, and that every action is available in every state, and that there are no independence assumptions between any states, actions, and transition probabilities. In other words, this is the result with no assumptions about the model structure. In practice, for an environment where only a small constant number of actions and transitions are available from a given state, many of the terms disappear.
Appendix B Additional Details on The Analysis of The RA-E3 Algorithm
@@
In this section, we comment on some additional details of the analysis of Theorem 13, the main theorem for the RA-E3 algorithm.
@@
We begin by defining approximation of MOMDPs. Think of $M$ as the true MOMDP, that is, a perfect model of transition probabilities. Think of $M^{\prime}$, on the other hand, as the current best estimate of an MOMDP obtained through exploration. In particular, $M^{\prime}$ in the RA-E3 algorithm will consist of the set of known states.
@@
(Kearns and Singh 2002) Let $M$ and $M^{\prime}$ be two MOMDPs over the same state space $\mathcal{S}$ with the same deterministic reward function $\mathbf{R}(s,a)$. $M^{\prime}$ is an $\alpha-$approximation of $M$ if for any state $s$ and $s^{\prime}$ and any action $a$,
@@
We now extend the Simulation Lemma (Kearns and Singh 2002), which tells us how close the approximation of an MOMDP needs to be in order for the expected welfare, or ESR, of a policy to be close in an estimated model. The argument is similar to (Kearns and Singh 2002), but in our case the policy may not be stationary and we want to bound the deviation in expected welfare rather than just accumulated reward.
@@
(Kearns and Singh 2002) Let $M$ be an MOMDP. A state $s$ of $M$ is considered known if it has be visited a number of times equal to
@@
(Kearns and Singh 2002) Let $M$ be an MOMDP. Let $s$ be a state of $M$ that has been visited at least $m$ times, with each action having been executed at least $\lfloor m/|\mathcal{A}|\rfloor$ times. Let $\hat{Pr}(s^{\prime}\mid s,a)$ denote the empirical transition probability estimates obtained from the $m$ visits to $s$. Then if
@@
We have specified the degree of approximation required for sufficient simulation accuracy. It remains to directly apply the Explore or Exploit Lemma from (Kearns and Singh 2002) to conclude the analysis.
Lemma 18 (Explore or Exploit Lemma (Kearns and Singh 2002)).
@@
Note that unlike the original E3 algorithm, which uses standard value iteration to compute the exactly optimal policy (optimizing a linear function of a scalar reward) on the sub-model $\hat{M}_{S}$, we use our RA-Value Iteration algorithm to find an approximately optimal policy. Therefore, we need to allocate $\epsilon/2$ error to both the simulation stage and the exploitation stage, which gives a total of $\epsilon$ error for the entire algorithm.
It remains to handle the failure parameter $\beta$ in the statement of the main theorem, which can be done similarly to (Kearns and Singh 2002). There are two sources of failure for the algorithm:
@@
Thus, the total computation time is bounded by $O\left(|\mathcal{S}||\mathcal{A}|(T/\alpha_{T\epsilon/2})^{d}\right)$ (the time required for RA-Value Iteration during off-line computations with $\epsilon/2$ precision by Lemma 11) times $T$ times the maximum number of attempted explorations, giving
$O((|\mathcal{S}|^{2}|\mathcal{A}|TG^{T}_{\max}/\epsilon)(T/\alpha_{T\epsilon/2})^{d}\log(|\mathcal{S}|/\beta)m_{known}).$
This concludes the remaining details of the analysis of the RA-E3 algorithm.

### v3

@@ Appendix A Proofs of Technical Lemmas
Appendix B Ablation Studies on Accumulated Rewards Discretization
We conduct further studies on the empirical effects of the discretization factor $\alpha$. For $W_{\text{SPF}}$, we set $\lambda=1$ (the smoothing factor). We gradually increase the disctretization interval of accumulated reward $\alpha$ gradually from $0.4$ to $1.2$ for a reward discount $\gamma=0.999$ for Taxi and $\gamma=0.99$ for Scavenger Hunt with the same settings as the experiments. As shown in Figure 4(a)(a) and Figure 4(b), we observe that the solution quality is often very high (better than baselines considered in the main body) even for discretization values that are substantially greater than those needed for theoretical guarantees (with $\epsilon=1$, we need $\alpha=2.5\times 10^{-3}$). Nonetheless the results are not necessarily monotone with respect to $\alpha$ in this regime.
(a) Nash Welfare in Taxi
(b) Cobb-Douglas in Scavenger Hunt
[…]
Figure 4: Ablation study on discretization factor $\alpha$.
This set of ablation studies demonstrates that the RAVI algorithm is capable of finding high-quality policies for optimizing expected scalarized return despite using a larger discretization parameter than that needed for our theoretical analysis. The practical takeaway is that, if one cannot afford the runtime for a very fine $\alpha$ with worst-case theoretical guarantees, one should instead try several much larger $\alpha$ values but still less than the smoothness of the objective. Due to the superlinear runtime scaling with respect to $\alpha$ for $d>1$, this can be substantially faster than a single run with much smaller $\alpha$.
Appendix C Details on the RAEE Algorithm
In this section, we give a detailed description of the RAEE algorithm, extending from [16] and demonstrating Theorem 2, the main theorem for the RAEE algorithm.
C.1 Introducing RAEE
@@
The extension works as follows. The algorithm starts off by doing balanced wandering [16]: when encountering a new state, the algorithm selects a random action. However, when revisiting a previously visited state, it chooses the least attempted action from that state, resolving ties by random action selection. At each state-action pair $(s,a)$ it tries, the algorithm stores the reward $\mathbf{R}(s,a)$ received and an estimate of the transition probabilities $Pr(s^{\prime}\mid s,a)$ derived from the empirical distribution of next states reached during balanced wandering.
@@
Next, we introduce the notion of a known state [16], which refers to a state that the algorithm has explored to the extent that the estimated transition probabilities for any action from that state closely approximate their actual values. Denote the number of times a state needs to be visited as $m_{known}$. We will specify the value later in our runtime characterization.
@@
States are thus categorized into three groups: known states, which the algorithm has extensively visited and obtained reliable transition statistics; states that have been visited before but remain unknown due to limited trials and therefore unreliable data; and states that have not been explored at all. By the Pigeonhole Principle, accurate statistics will eventually accumulate in some states over time, leading to their becoming known. Let $S$ be the set of currently known states, the algorithm can build the current known-state MOMDP $M_{S}$ that is naturally induced on $S$ by the full MOMDP $M$ with all “unknown” states merged into a single absorbing state. Although the algorithm cannot access $M_{S}$ directly, it will have an approximation $\hat{M}_{S}$ by the definition of the known states. By the simulation lemma [16], $\hat{M}_{S}$ will be an accurate model in the sense that the expected $T$-step welfare of any policy in $\hat{M}_{S}$ is close to its expected $T$-step return in $M_{S}$. (Here $T$ is the horizon time.) Hence, at any timestep, $\hat{M}_{S}$ functions as an incomplete representation of $M$, specifically focusing on the part of $M$ that the algorithm possesses a strong understanding of.
@@
This is where we insert RAVI. The algorithm performs the two off-line optimal policy computations; i) first on $\hat{M}_{S}$ using RAVI to compute an exploitation policy that yields an approximately optimal welfare and ii) second performing traditional value iteration on $\hat{M}^{\prime}_{S}$, which has the same transition probabilities as $\hat{M}_{S}$, but different payoffs: in $\hat{M}^{\prime}_{S}$, the absorbing state (representing “unknown” states) has scalar reward $1$ and all other states have scalar reward 00. The optimal policy in $\hat{M}^{\prime}_{S}$ simply exits the known model as rapidly as possible, rewarding exploration.
@@
By the explore or exploit lemma [16], the algorithm is guaranteed to either output a policy with approximately optimal return in $M$, or to improve the statistics at an unknown state. Again by the Pigeonhole Principle, a new state becomes known after the latter case occurs for some finite number of times, and thus the algorithm is always making progress. In the worst case, the algorithm builds a model of the entire MOMDP $M$. Having described the elements, we now outline the entire extended algorithm, where the notations are consistent with [16].
RAEE Algorithm:
@@
  • (Discovery of New Known States) Any time a state $s$ has been visited $m_{known}$ [16] times during balanced wandering, it enters the known set $S$, and no longer participates in balanced wandering.
@@
    • (Attempted Exploitation) Use RAVI algorithm to compute an $\epsilon/2$-optimal policy on $\hat{M}_{S}$. If the resulting exploitation policy $\hat{\pi}$ achieves return from $s$ in $\hat{M}_{S}$ that is at least $V^{*}(s,\mathbf{0},T)-\epsilon/2$, the algorithm halts and outputs $\hat{\pi}$.
@@
This concludes the description of the algorithm.
C.2 Runtime Analysis
@@
In this subsection, we comment on additional details of the analysis of Theorem 2, the main theorem for the RAEE algorithm.
Theorem 2 (RAEE). Let $V^{*}(s,0,T)$ denote the value function for the policy with the optimal expected welfare in the MOMDP $M$ starting at state $s$, with $\mathbf{0}\in\mathbb{R}^{d}$ accumulated reward and $T$ timesteps remaining. Then for a uniformly continuous welfare function $W$, there exists an algorithm $A$, taking inputs $\epsilon$, $\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, and $V^{*}(s,\mathbf{0},T)$, such that the total number of actions and computation time taken by $A$ is polynomial in $1/\epsilon$, $1/\beta$, $|\mathcal{S}|$, $|\mathcal{A}|$, the horizon time $T=1/(1-\gamma)$ and exponential in the number of objectives $d$, and with probability at least $1-\beta$, $A$ will halt in a state $s$, and output a policy $\hat{\pi}$, such that $V^{\hat{\pi}}_{M}(s,0,T)\geq V^{*}(s,0,T)-\epsilon$.
@@
We begin by defining approximation of MOMDPs. Think of $M$ as the true MOMDP, that is, a perfect model of transition probabilities. Think of $M^{\prime}$, on the other hand, as the current best estimate of an MOMDP obtained through exploration. In particular, $M^{\prime}$ in the RAEE algorithm will consist of the set of known states.
@@
[16] Let $M$ and $M^{\prime}$ be two MOMDPs over the same state space $\mathcal{S}$ with the same deterministic reward function $\mathbf{R}(s,a)$. $M^{\prime}$ is an $\alpha-$approximation of $M$ if for any state $s$ and $s^{\prime}$ and any action $a$,
@@
We now extend the Simulation Lemma [16], which tells us how close the approximation of an MOMDP needs to be in order for the expected welfare, or ESR, of a policy to be close in an estimated model. The argument is similar to [16], but in our case the policy may not be stationary and we want to bound the deviation in expected welfare rather than just accumulated reward.
@@
[16] Let $M$ be an MOMDP. A state $s$ of $M$ is considered known if it has be visited a number of times equal to
@@
[16] Let $M$ be an MOMDP. Let $s$ be a state of $M$ that has been visited at least $m$ times, with each action having been executed at least $\lfloor m/|\mathcal{A}|\rfloor$ times. Let $\hat{Pr}(s^{\prime}\mid s,a)$ denote the empirical transition probability estimates obtained from the $m$ visits to $s$. Then if
@@
We have specified the degree of approximation required for sufficient simulation accuracy. It remains to directly apply the Explore or Exploit Lemma from [16] to conclude the analysis.
Lemma 7 (Explore or Exploit Lemma [16]).
@@
Note that unlike the original E3 algorithm, which uses standard value iteration to compute the exactly optimal policy (optimizing a linear function of a scalar reward) on the sub-model $\hat{M}_{S}$, we use our RAVI algorithm to find an approximately optimal policy. Therefore, we need to allocate $\epsilon/2$ error to both the simulation stage and the exploitation stage, which gives a total of $\epsilon$ error for the entire algorithm.
It remains to handle the failure parameter $\beta$ in the statement of the main theorem, which can be done similarly to [16]. There are two sources of failure for the algorithm:
@@
Thus, the total computation time is bounded by $O\left(|\mathcal{S}|^{2}|\mathcal{A}|(T/\alpha_{T\epsilon/2})^{d}\right)$ (the time required for RAVI during off-line computations with $\epsilon/2$ precision by Lemma 3) times $T$ times the maximum number of attempted explorations, giving
$O((|\mathcal{S}|^{3}|\mathcal{A}|TG^{T}_{\max}/\epsilon)(T/\alpha_{T\epsilon/2})^{d}\log(|\mathcal{S}|/\beta)m_{known}).$
This concludes the proof of Theorem 2.